OpenStudy (anonymous):

suppose the population of deer in a state is 10,280 and growing 1% each year. predict the population after five years.

jimthompson5910 (jim_thompson5910):

Your equation would be \[\Large P = 10280(1+0.01)^t\] \[\Large P = 10280(1.01)^t\] plug in t = 5 and evaluate to find P